# Confidential Computing on nVIDIA H100 GPU: A Performance Benchmark Study

### v2

@@ Abstract
This report evaluates the performance impact of enabling Trusted Execution Environments (TEE) on nVIDIA H100 GPUs for large language model (LLM) inference tasks. We benchmark the overhead introduced by TEE mode across various LLMs and token lengths, with a particular focus on the bottleneck caused by CPU-GPU data transfers via PCIe. Our results indicate that while there is minimal computational overhead within the GPU, the overall performance penalty is primarily attributable to data transfer. For the majority of typical LLM queries, the overhead remains below 5%, with larger models and longer sequences experiencing nearly zero overhead.
@@ 1 Introduction
Trusted Execution Environments (TEEs) are increasingly important in machine learning and AI due to growing security requirements in both enterprise and decentralized applications [SAB15, MSM+18, AKKH18]. The introduction of TEE-enabled GPUs, such as the nVIDIA H100, adds an extra layer of protection for sensitive data but may impact performance. Understanding these trade-offs, particularly for large-scale machine learning tasks, is crucial for adopting TEE in high-performance AI applications [YMY+22, WO24].
This report quantifies the performance overhead of enabling TEE on the nVIDIA H100 GPU during LLM inference tasks, identifying where the overhead arises and under what conditions it can be minimized.
@@ 2.1 Trusted Execution Environment
2.2 nVIDIA H100 GPU
The nVIDIA H100 GPU marks a significant milestone as the first GPU to support TEE. In TEE mode, the H100 operates in an isolated and secure environment where data transfers between the CPU and GPU are encrypted. This is achieved through “bounce buffers”, which protect all inputs and outputs during transit between the CPU’s encrypted memory and the GPU’s internal memory [DGK+23].
To maintain end-to-end security, the H100 works in conjunction with CPU TEEs, such as Intel’s TDX [Int] or AMD’s SEV-SNP [AMD, SS20], securing communication channels between the GPU driver and interacting software. This setup prevents unauthorized access and ensures data integrity throughout the process.
The H100 also implements remote attestation to verify the GPU’s identity and the authenticity of its firmware. Additionally, Secure Boot ensures that only authenticated firmware is executed during the GPU’s boot process, further strengthening security.
@@ 2.3 Performance Impact
Enabling TEE on the nVIDIA H100 GPU introduces performance overheads primarily due to additional encryption and decryption during secure data transfer [MYF+24]. While the GPU’s internal computation remains unaffected, the main bottleneck lies in the CPU-GPU I/O, particularly when data is exchanged via PCIe. This impact varies with the size of the data transfer. The following sections present experimental results quantifying these effects across various use cases.
With the TEE-enabled nVIDIA H100 GPU, it becomes crucial to quantify performance trade-offs during practical use cases. In the next section, we outline the methodology used to assess the performance impact of TEE during LLM inference tasks.
@@ 3.2 Test Scenarios
  • TEE mode ON vs. TEE mode OFF: Tests were performed with TEE mode alternately enabled and disabled on the H100 GPU, allowing for a direct comparison of performance.
@@ 3.3.1 Infrastructure
The experiments were set up with the following hardware:
  • GPU: nVIDIA H100 NVL (94GB, 3.9TB/s bandwidth)
  • CPU: AMD EPYC 9V84 96-Core Processor with SEV-SNP
  • Memory: 314 GB
  • Driver versions:
    • CUDA 12.5 (driver version 555.42.02)
    • Kernel driver version 550.90.07
@@ 3.3.3 Models
  • Meta-Llama-3.1-70B-Instruct with 4-bit bitsandbytes quantization to fit into a single H100 GPU
@@ 4 Results
Conclusion 1: The average overhead is less than 7%. We quantified the overhead by measuring the throughput with TEE mode enabled versus disabled, across varying input sizes and model configurations, as shown in Table 1.
Model TPS (tokens/s) QPS (req/s)
@@
Conclusion 2: The overhead reduces toward zero as the model size grows. As shown in Table 1, the smallest model (Llama-3.1-8B) has the highest overhead. The medium-sized model (Phi-3-14B-128k) has roughly two-thirds of the overhead compared to the smaller one. The largest model (Llama-3.1-70B) has a trivial overhead close to zero.
Conclusion 3: The latency is the main factor contributing to the overhead of the TEE mode. Table 2 shows the overhead introduced to the latency measured by TTFT and ITL. TTFT has a higher overhead compared with ITL, indicating the bottleneck is likely introduced by the I/O instead of the computation happening inside the TEE. Nevertheless, the overhead becomes trivial when hosting heavy computation models like Llama-3.1-70B.
Model TTFT (s) ITL (s)
@@
Model TPS - short (tokens/s) TPS - medium (tokens/s) TPS - long (tokens/s)
@@
Conclusion 5: TEE can reach typical throughput
Our experiments revealed that, with medium-sized inputs, the nVIDIA H100 GPU achieves 130 TPS for Llama-3.1-8B, while the larger Phi-3-14B model reaches approximately 6 TPS. These results demonstrate the robust performance of the H100 GPU across models of varying complexity.
More detailed experimental data is shown in Figure 2, 3, and 4.
Figure 2: Throughput vs output token size for LLama-3.1-8B
Figure 3: Throughput vs output token size for Phi3-14B-128k
Figure 4: Throughput vs output token size for Llama-3.1-70B

### v3

# Confidential Computing on nVIDIA Hopper GPUs: A Performance Benchmark Study

@@ Abstract
This report evaluates the performance impact of enabling Trusted Execution Environments (TEE) on nVIDIA Hopper GPUs for large language model (LLM) inference tasks. We benchmark the overhead introduced by TEE mode across various LLMs and token lengths, with a particular focus on the bottleneck caused by CPU-GPU data transfers via PCIe. Our results indicate that while there is minimal computational overhead within the GPU, the overall performance penalty is primarily attributable to data transfer. For the majority of typical LLM queries, the overhead remains below 7%, with larger models and longer sequences experiencing nearly zero overhead.
Acknowledgments
We would like to express our gratitude to the io.net [io.] and IOG Foundation [Fou] for their generous grant, which made this research possible. We also extend our thanks to Engage Stack [Sta], the cloud service provider, for providing the necessary hardware and technical support.
@@ 1 Introduction
Trusted Execution Environments (TEEs) are increasingly important in machine learning and AI due to growing security requirements in both enterprise and decentralized applications [SAB15, MSM+18, AKKH18]. The introduction of TEE-enabled GPUs, such as the nVIDIA H100 and H200, adds an extra layer of protection for sensitive data but may impact performance. Understanding these trade-offs, particularly for large-scale machine learning tasks, is crucial for adopting TEE in high-performance AI applications [YMY+22, WO24].
This report quantifies the performance overhead of enabling TEE on the nVIDIA Hopper architecture GPUs during LLM inference tasks, identifying where the overhead arises and under what conditions it can be minimized.
@@ 2.1 Trusted Execution Environment
2.2 nVIDIA Hopper Architecture
The nVIDIA Hopper architecture marks a significant milestone as the first GPU family to support TEE [nVI]. In TEE mode, the GPU operates in an isolated and secure environment where data transfers between the CPU and GPU are encrypted. This is achieved through “bounce buffers”, which protect all inputs and outputs during transit between the CPU’s encrypted memory and the GPU’s internal memory [DGK+23].
To maintain end-to-end security, the GPU works in conjunction with CPU TEEs, such as Intel’s TDX [Int] or AMD’s SEV-SNP [AMD, SS20], securing communication channels between the GPU driver and interacting software. This setup prevents unauthorized access and ensures data integrity throughout the process.
The Hopper series also implement remote attestation to verify the GPU’s identity and the authenticity of its firmware. Additionally, Secure Boot ensures that only authenticated firmware is executed during the GPU’s boot process, further strengthening security.
@@ 2.3 Performance Impact
Enabling TEE on the nVIDIA Hopper GPU introduces performance overheads primarily due to additional encryption and decryption during secure data transfer [MYF+24]. While the GPU’s internal computation remains unaffected, the main bottleneck lies in the CPU-GPU I/O, particularly when data is exchanged via PCIe. This impact varies with the size of the data transfer. The following sections present experimental results quantifying these effects across various use cases.
With the TEE-enabled nVIDIA Hopper GPU, it becomes crucial to quantify performance trade-offs during practical use cases. In the next section, we outline the methodology used to assess the performance impact of TEE during LLM inference tasks.
@@ 3.2 Test Scenarios
  • TEE mode ON vs. TEE mode OFF: Tests were performed with TEE mode alternately enabled and disabled on the H-series GPUs, allowing for a direct comparison of performance.
@@ 3.3.1 Infrastructure
The experiments were set up with the following hardwares, respectively.
Component Specification Setup 1 Setup 2
GPU Model NVIDIA H100 NVL NVIDIA H200 NVL
Memory 94 GB 141 GB
Bandwidth 3.9 TB/s 4.8 TB/s
CPU Model AMD EPYC 9V84 INTEL XEON PLATINUM 8558
Cores 96 48
TEE Technology SEV-SNP TDX
Memory Total Memory 314 GB 128 GB
Software CUDA Version 12.5 12.5
Driver Version 555.42.06 555.42.06
Kernel Driver Version 550.90.07 550.90.07
Table 1: Hardware Setup
@@ 3.3.3 Models
  • Meta-Llama-3.1-70B-Instruct with 4-bit bitsandbytes quantization to fit into a single Hopper GPU
@@ 4 Results
Conclusion 1: The average overhead is less than 9%. We quantified the overhead by measuring the throughput with TEE mode enabled versus disabled, across varying input sizes and model configurations, as shown in Table 2.
GPU Model TPS (tokens/s) QPS (req/s)
@@
We observed a difference in the impact of TEE mode between H100 and H200 in Table 2. First, compared to H100, TEE mode introduces more overhead in H200 when runing the same model. Second, the impact of TEE mode on QPS is more noticeable in H200. In H100, the impact of TEE mode on TPS is nearly double that of QPS, but in H200, the difference decreased significantly.
Conclusion 2: The overhead reduces as the model size grows. As shown in Table 2, the smallest model (Llama-3.1-8B) has the highest overhead. The medium-sized model (Phi-3-14B-128k) has roughly two-thirds of the overhead compared to the smaller one. Notably, the largest model (Llama-3.1-70B) has a negligible overhead close to zero in H100.
Conclusion 3: The latency is the main factor contributing to the overhead of the TEE mode. Table 3 shows the overhead introduced to the latency measured by TTFT and ITL. TTFT has a higher overhead compared with ITL, indicating the bottleneck is likely introduced by the I/O instead of the computation happening inside the TEE. Nevertheless, the overhead becomes trivial when hosting heavy computation models like Llama-3.1-70B. Additionally, TEE mode has also a greater impact on TTFL and ITL in H200 compared to H100.
GPU Model TTFT (s) ITL (s)
@@
GPU Model TPS - short (tokens/s) TPS - medium (tokens/s) TPS - long (tokens/s)
@@
Conclusion 5: TEE can reach typical throughput. Here, we use nVIDIA H100 for the case study. Our experiments revealed that, with medium-sized inputs, the H100 GPU achieves 130 TPS for Llama-3.1-8B, while the larger Phi-3-14B model reaches approximately 6 TPS. These results demonstrate the robust performance of the H100 GPU across models of varying complexity.
More detailed experimental data for H100 is shown in Figures 2, 3, and 4, and for H200 in Figures 5, 6, and 7.
Figure 2: Throughput vs output token size for LLama-3.1-8B in H100.
Figure 3: Throughput vs output token size for Phi3-14B-128k in H100.
Figure 4: Throughput vs output token size for Llama-3.1-70B in H100.
Figure 5: Throughput vs output token size for LLama-3.1-8B in H200.
Figure 6: Throughput vs output token size for Phi3-14B-128k in H200.
Figure 7: Throughput vs output token size for Llama-3.1-70B in H200.